# Differentiable Voxelization and Mesh Morphing

### v1

@@ Abstract.
This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
@@ 1. Introduction
Figure 1. Differentiable voxelization algorithm converts a 3D mesh into a 3D grid of occupancy voxels by the winding number and solid angles on GPU. The first row shows hard cases of the voxelization of 3D meshes using traditional methods, where the mesh is not watertight, or containing non-regular triangles, or formed by combinations of shapes, or with complex geometry. The second row shows the voxelization results of the proposed differentiable voxelization algorithm followed by the re-cubifying.
@@
Voxel is a fundamental representation of 3D objects in computer graphics and computer vision. It is widely used in 3D reconstruction (Brock et al., 2016; Wang et al., 2022), 3D rendering (Kato et al., 2020; Loper and Black, 2014; Liu et al., 2019), and 3D generative models (Mescheder et al., 2019; Guo et al., 2023; Ren et al., 2024) due to its uniform grid structure and simplicity. Therefore, voxelization, which converts a 3D mesh into a 3D grid of voxels, is an essential preprocessing step for many 3D deep learning tasks. However, the commonly used voxelization algorithms, such as the rasterization-based and the mesh-boolean-based methods (Wang, 2010; Zhou et al., 2016; Wang et al., 2022), implemented on the CPU, are not efficient and accurate, which comsumes a lot of time and memory and limits the development of high-resolution 3D deep learning works. To address this issue, we propose a novel differentiable voxelization algorithm for 3D meshes, which is fast by GPU acceleration, flexible in terms of abitrary smapling resolution on non-watertight meshes with abitrary topologies. More importantly, the proposed algorithm is differentiable, which enables the computation of gradients with respect to the input mesh, facilitating the optimization of mesh-level deep learning tasks. The proposed approach is based on the winding number and solid angles.
@@
In recent work, Luo et al. (Luo et al., 2022) used the abstract inverse quadratic field to compute the occupancy (Mescheder et al., 2019) of voxels and slice a mesh differentiablely, which facilitates the slicing and voxelization of 3D meshes and achieves mesh reconstruction from 2D medical scan images. However, for general 3D meshes with non-regualr topologies and wiring structures, the abstract inverse quadratic field is not accurate enough due to the integration error. This issue is addressed by replacing field integration with the solid angles, which are more accurate and efficient in computing the occupancy of voxels without loss of differentiability. Jacobson et al. (Jacobson et al., 2013; Jacobson, 2016) achieved the voxelization and boolean operations of 3D meshes by the winding number and solid angles, but the naive computation of the solid angles leads a computational gap near surface boundaries, which requires further human interventions for discretization and smoothing. Further more, we discard the hierarchical winding number evaluation from (Jacobson et al., 2013) but implement the solid angles computation in a parallel manner on GPU, which is more efficient and accurate. Leveraging the proposed differentiable voxelization, we further demonstrate the application of the proposed voxelization in mesh morphing.
@@
  • It is a fast and resolution-flexible algorithm for voxelization of 3D meshes, addmiting GPU-accelerated.
  • It is a differentiable algorithm for voxelization of 3D meshes, facilitating mesh morphing and other mesh-level optimization and deep learning tasks.
  • The proposed algorithm achieves state-of-the-art performance in terms of both accuracy and efficiency on the ShapeNet dataset.
@@ 2.1. Generalized Winding Number
$\vec{E_{q}}$ $=\frac{\vec{r}-\vec{p}}{\left\|\vec{r}-\vec{p}\right\|^{3}}.$
For a closed boundary surface $\Sigma=\partial\Omega$, the winding number around the query point $p$ is equivalent the occupancy, which is a direct consequence of the Gaussian integral theorem (Maxwell, 1954), i.e.,
@@
(4) $W(p,\Sigma)=$ $\frac{1}{4\pi}\sum_{i=1}^{F}\left\langle\vec{n}_{i},\vec{E_{q}}(c_{i})\right\rangle\Delta_{i},$
@@ 2.2. Solid Angles
However, general 3D meshes, especially the artifact meshes from manually designing and manufacturing, usually contain sharp edges, non-regular triangles, which leads to considerable errors in the naive discretization of the winding number. To address this issue, (Jacobson et al., 2013) proposed the hierarchical winding number evaluation by calulating solid angles, i.e.,
(5) $\text{Occp}(v,\Sigma)=W(p,\Sigma)=$ $\frac{1}{4\pi}\sum_{i=1}^{F}\Omega_{i}(q),$
@@
The naive computation of the solid angles like Eq. (5) leads a computational gap near surface boundaries due to the zero point of the denominator. (Luo et al., 2022) involves discretization and smoothing to achieve binary occupancy, which declines the accuracy. We replace $arctan$ by $\arctan 2$ to address this issue, where we prove that the $\arctan 2$ is continuous and smooth with zero gradients over query points coordinates, except the query points on the surface. Refer to the supplementary material for the proof. The proposed differentiable voxelization algorithm is presented by Algorithm 1.
@@
Fig. 1 shows an illustration of the differentiable voxelization of 3D meshes, including non-watertight meshes, meshes with non-regular triangles, and combinational and complex geometry, which causes the failure of traditional voxelization methods. The proposed differentiable voxelization algorithm achieves accurate and efficient voxelization of 3D meshes, which is essential for mesh morphing and other mesh-level optimization and deep learning tasks.
@@ 2.3. Mesh Morphing
Differentiable voxelization facilitates the optimization of mesh morphing, which admits the computation of gradients with respect to the input mesh under the supervision or conditions of voxelization ground truth. The voxelization ground truth can be usually obtained by CT scans (Buzug, 2011). Therefore the differentiable voxelization lays the foundation for 3D mesh reconstruction from 2D medical scan images (Litjens et al., 2017). It is worth noting that the differentiable voxelization is not limited to mesh morphing, but can be incorporated in other mesh-level optimization and deep learning tasks, such as mesh segmentation, mesh classification, and mesh generation.
Algorithm 1 provided an accurate differentiable voxelization algorithm for mesh morphing, which means the Eq. (5) converges to the ground truth of the voxelization. Even though Eq. (5) is composed by differentiable functions, the results approach the binary values of the voxelization ground truth. However, pros always come with cons. High accuracy of binary voxelization leads to gradients vanishing, despit the calculation is differentiable. In practice of mesh reconstruction of medical images, we use (4) to avoid the vanishing gradients, which is a trade-off between the accuracy and the differentiability. Fig 2 shows an illustration of the mesh morphing to fit the target segmentation of a human heart. Refer to (Luo et al., 2022) for more details of the cardiac mesh reconstruction via the differentiable voxelization.
@@ 3. Experiments
In the experiments, we evaluate our differentiable voxelization algorithm and other exiting methods on some of the meshes from (Krishnamurthy and Levoy, 1996) and (Jacobson, 2022). The methods evaluated includes the ”contains” method in Trimesh (Dawson-Haggerty et al., [n.d.]) library, the Mesh-to-SDF library (Kleineberg, 2024), and the SDF computing method used in (Wang et al., 2022), where we denoted as DOGN.
We evaluate the methods under different resolutions from 32 to 256, (32, 64, 128, 256). The Mesh-to-SDF method has two kinds of surface point methods namely ’scan’ and ’sample’. The ’Scan’ method, DOGN and our method are evaluated under all resolutions from 32 to 256. The ’Sample’ method and ’contains’ method are only evaluated under 32 and 64 which we think are sufficient to demonstrate their performance in terms of quality and time.
@@ 3.1. Settings and Details of the Methods
    Ours: since our method relies on GPU to process the data in parallel, we set the batch size of our method, i.e. the number of query points of each batch, to 2000, due to the memory limitation. Also, all the calculation is done under half-precision.
@@
    Mesh-to-sdf (scan): For ’scan’ we used ’mesh-to-voxels’ API and set different resolutions directly.
@@
All the meshes are normalized to [-1,1] before the computation of the ccupancy filed (SDF). All occupancy fields are firstly cubified and then smoothed with filter Laplacian method in Trimesh with 10 iterations and $\lambda=0.15$. Then, we compute the Chamfer distance and the Hausdorff distance with 20000 samples between the normalized original mesh and smoothed mesh 3 times, with the standard deviation reported. Results are computed under a single Nvidia RTX A6000 GPU and an AMD EPYC 7443P 24-core Processor.
@@ 3.2. Results
It can be seen from table 1 that our method significantly outperforms ’Scan’ and DOGN in terms of quality (for results with standard deviation please refer to table 6 in Appendix A), under the resolution of 256. The cases in which the two distances have a large difference, the Max Planck for instance, is because the original mesh is not closed, and our method tends to fill those holes. The resulting mesh still captures the details of the original mesh and looks similar, as shown in figure give a car picture here ! . The time cost of our method is closely related to the number of faces of the mesh and the memory of the GPU, since our method relies on GPU memory, and computes the integral over all faces of a mesh. For example, our method processes mesh ’Ladder’ and ’Suzanne’ faster than the other two methods, but is significantly slower then process complex meshes like ’Igea’. However, with sufficient computing resources, our method will be significantly faster as shown in table 2. Otherwise, the processing time grows almost linearly to the number of faces and cubicly to the resolution when computing resources are limited.

### v2

@@ Abstract.
This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing yihaoluo2020@gmail.com. Copyright is held by the owner/author(s).
* Corresponding author.
$\dagger$ Co-last Senior authors.
@@ 1. Introduction
Figure 1. Differentiable voxelization algorithm converts a 3D mesh into a 3D grid of occupancy voxels by the winding number and solid angles on GPU. The first row shows hard cases of the voxelization of 3D meshes using traditional methods, where the mesh is not watertight, contains irregular triangles, or is formed by combinations of shapes or with complex geometry. The second row shows the voxelization results of the proposed differentiable voxelization algorithm, followed by the re-cubifying.
@@
Voxel is a fundamental representation of 3D objects in computer graphics and computer vision. It is widely used in 3D reconstruction (Brock et al., 2016; Wang et al., 2022), 3D rendering (Kato et al., 2020; Loper and Black, 2014; Liu et al., 2019), and 3D generative models (Mescheder et al., 2019; Guo et al., 2023; Ren et al., 2024) due to its uniform grid structure and simplicity. Therefore, voxelization, which converts a 3D mesh into a 3D grid of voxels, is an essential preprocessing step for many 3D deep learning tasks. However, the commonly used voxelization algorithms, such as the rasterization-based and the mesh-boolean-based methods (Wang, 2010; Zhou et al., 2016; Wang et al., 2022), implemented on the CPU, are not efficient and accurate, which consumes a lot of time and memory and limits the development of high-resolution 3D deep learning works. To address this issue, we propose a novel differentiable voxelization algorithm for 3D meshes, which is fast by GPU acceleration and flexible in terms of arbitrary sampling resolution on non-watertight meshes with arbitrary topologies. More importantly, the proposed algorithm is differentiable, which enables the computation of gradients with respect to the input mesh, facilitating the optimization of mesh-level deep learning tasks. The proposed approach is based on the winding number and solid angles.
@@
In recent work, Luo et al. (Luo et al., 2022) used the abstract inverse quadratic field to compute the occupancy (Mescheder et al., 2019) of voxels and slice a mesh differentiablely, which facilitates the slicing and voxelization of 3D meshes and achieves mesh reconstruction from 2D medical scan images. However, for general 3D meshes with irregular topologies and wiring structures, the abstract inverse quadratic field is not accurate enough due to the integration error. This issue is addressed by replacing field integration with solid angles, which are more accurate and efficient in computing the occupancy of voxels without loss of differentiability. Jacobson et al. (Jacobson et al., 2013; Jacobson, 2016) achieved the voxelization and boolean operations of 3D meshes by the winding number and solid angles, but the naive computation of the solid angles leads to a computational gap near surface boundaries, which requires further human interventions for discretization and smoothing. Furthermore, we discard the hierarchical winding number evaluation from (Jacobson et al., 2013) but implement the solid angles computation in a parallel manner on GPU, which is more efficient and accurate. Leveraging the proposed differentiable voxelization, we further demonstrate the application of the proposed voxelization in mesh morphing.
@@
  • It is a fast and resolution-flexible algorithm for the voxelization of 3D meshes endowed with GPU acceleration.
  • It is a differentiable algorithm for the voxelization of 3D meshes, facilitating mesh morphing and other mesh-level optimization and deep learning tasks.
  • The proposed algorithm achieves state-of-the-art performance on the mesh-to-voxel task in terms of both accuracy and efficiency on the ShapeNet dataset.
@@ 2.1. Generalized Winding Number
$\vec{E_{q}}(x)$ $=\frac{\vec{x}-\vec{q}}{\left\|\vec{x}-\vec{q}\right\|^{3}}.$
For a closed boundary surface $\Sigma=\partial\Omega$, the winding number around the query point $p$ is equivalent to the occupancy, which is a direct consequence of the Gaussian integral theorem (Maxwell, 1954), i.e.,
@@
(4) $W(q,\Sigma)=$ $\frac{1}{4\pi}\sum_{i=1}^{F}\left\langle\vec{n}_{i},\vec{E_{q}}(c_{i})\right\rangle\Delta_{i},$
@@ 2.2. Solid Angles
However, general 3D meshes, especially the artifact meshes from manually designing and manufacturing, usually contain sharp edges and irregular triangles, which leads to considerable errors in the naive discretization of the winding number. To address this issue, (Jacobson et al., 2013) proposed the hierarchical winding number evaluation by calculating solid angles, i.e.,
(5) $\text{Occp}(q,\Sigma)=W(p,\Sigma)=$ $\frac{1}{4\pi}\sum_{i=1}^{F}\Omega_{i}(q),$
@@
[…]
@@
Fig. 1 shows an illustration of the differentiable voxelization of 3D meshes, including non-watertight meshes (almost-closed meshes), meshes with non-regular triangles, and combinational and complex geometry, which causes the failure of traditional voxelization methods. More importantly, the proposed differentiable voxelization algorithm can be easily implemented on GPU, which achieves fast and accurate voxelization of 3D meshes.
2.3. Voxelization of Open Surfaces
Although the differentiable voxelization has no limits on non-watertight meshes by automatically filling the miss faces through continuous integration, as shown in Fig. 1 (a), it is still a challenge to voxelize completely open meshes, which is a common case for artist-designed models. To address this issue, we propose the flipped duplication to convert the open meshes into almost closed meshes, described as the following algorithm 2.
Algorithm 2 Flipping Duplication
0:  Open Mesh $\Sigma=(V,F)$
0:  Flipped Duplication $\Sigma^{-2}=(V\cup V^{\prime},F\cup F^{\prime})$
  Compute normal vectors $N_{i}$ for each vertices $v_{i}$ in $V$
  Copy the vertices and perturb them towards the negative normal direction, $v^{\prime}=v_{i}-\epsilon N_{i}$, $V^{\prime}=\{v^{\prime}|v_{i}\in V\}$ (where $\epsilon$ is a small positive number, e.g., $0.01$)
  Copy and flip the faces
  $F^{\prime}=\{(v_{0}^{\prime},v_{1}^{\prime},v_{2}^{\prime})|f=(v_{0},v_{1},v_{2})\in F\}$
  Return $\Sigma^{-2}=(V\cup V^{\prime},F\cup F^{\prime})$
Apply Algorithm 1 on the flipped duplication $\Sigma^{-2}$ obtained by Algorithm 2, the open meshes can be voxelized as the closed meshes, which is shown in Fig. 3. The thickness of the slit between the flipped duplication is determined by the perturbation $\epsilon$ in Algorithm 2. Algorithm 1 can easily overcome the non-watertight issue caused by the slit.
Figure 3. The flipped duplication converts the open meshes into almost closed meshes, which can be voxelized by the proposed differentiable voxelization algorithm.
@@ 2.4. Mesh Morphing
Differentiable voxelization facilitates the optimization of mesh morphing, which admits the computation of gradients with respect to the input mesh under the supervision or conditions of voxelization ground truth. The voxelization ground truth can be usually obtained by CT scans (Buzug, 2011). Therefore the differentiable voxelization lays the foundation for 3D mesh reconstruction from 2D medical scan images (Litjens et al., 2017). It is worth noting that the differentiable voxelization is not limited to mesh morphing but can be incorporated in other mesh-level optimization and deep learning tasks, such as mesh segmentation, mesh classification, and mesh generation.
Algorithm 1 provided an accurate differentiable voxelization algorithm for mesh morphing, which means the Eq. (5) converges to the ground truth of the voxelization. Even though Eq. (5) is composed by differentiable functions, the results approach the binary values of the voxelization ground truth. However, pros always come with cons. High accuracy of binary voxelization leads to gradients vanishing, despite the calculation being differentiable. In the practice of mesh reconstruction of medical images, we use (4) to avoid the vanishing gradients, which is a trade-off between the accuracy and the differentiability. Fig 4 shows an illustration of the mesh morphing to fit the target segmentation of a human heart. Refer to (Luo et al., 2022) for more details of the cardiac mesh reconstruction via the differentiable voxelization.
@@ 3. Experiments
In the experiments, we evaluate our differentiable voxelization algorithm on the Mesh-to-occupancy task compared with other existing methods on selected meshes from (Krishnamurthy and Levoy, 1996) and (Jacobson, 2022). The methods evaluated include the ”contains” method in Trimesh (Dawson-Haggerty et al., [n.d.]) library, the Mesh-to-SDF library (Kleineberg, 2024), and the SDF computing method used in (Wang et al., 2022), where we denoted as DOGN.
Direct evaluation of the occupancy field is not available due to the lack of ground truth. In fact, a more accurate occupancy field will yield a better reconstruction mesh via marching cubes. We measure the geometric distance between the reconstructed and original mesh to evaluate the mesh-to-voxel algorithms. We use Hausofroff and Chamfer distances as the main metrics.
We evaluate the methods at different resolutions, from 32 to 256 (32, 64, 128, 256). The Mesh-to-SDF method has two kinds of surface point methods namely ’scan’ and ’sample’. The ’Scan’ method, DOGN and our method are evaluated under all resolutions from 32 to 256. The ’Sample’ method and ’contains’ method are only evaluated under 32 and 64 which we think are sufficient to demonstrate their performance in terms of quality and time.
@@ 3.1. Settings and Details of the Methods
    Ours: since our method relies on GPU to process the data in parallel, we set the batch size of our method, i.e., the number of query points of each batch, to 2000 due to the memory limitation. Also, all the calculations are done under half-precision.
@@
    Mesh-to-sdf (scan): For ’scan’, we used the ’mesh-to-voxels’ API and set different resolutions directly.
@@
All the meshes are normalized to [-1,1] before the computation of the occupancy filed (SDF). All occupancy fields are firstly cubified and then smoothed with filter Laplacian method in Trimesh with ten iterations and $\lambda=0.15$. Then, we compute the Chamfer distance and the Hausdorff distance with 20000 samples between the normalized original mesh and smoothed mesh 3 times, with the standard deviation reported. Results are computed under a single Nvidia RTX A6000 GPU and an AMD EPYC 7443P 24-core Processor.
@@ 3.2. Results
It can be seen from table 1 that our method significantly outperforms ’Scan’ and DOGN in terms of quality (for results with standard deviation, please refer to table 6 in Appendix A), under the resolution of 256. In cases where the two distances have a large difference, ’Car’ for instance, is because the original mesh is non-watertight, and our method fill those holes and splits. The resulting mesh still captures the details of the original mesh and looks similar, shown in Fig. 1 (c). The time cost of our method is closely related to the number of faces of the mesh and the memory of the GPU, since our method relies on GPU memory, and computes the integral over all faces of a mesh. For example, our method processes mesh ’Ladder’ and ’Suzanne’ faster than the other two methods, but is significantly slower then process complex meshes like ’Igea’. However, with sufficient computing resources, our method will be significantly faster as shown in table 2. Otherwise, the processing time grows almost linearly to the number of faces and cubicly to the resolution when computing resources are limited.